OpenStudy (anonymous):

help, please. Let f(x) = 2x2 – 2x + 7 and g(x) = x2 + 9x – 6. Find f(x) – g(x). x2 – 11x + 13 x2 – 11x + 1 x2 + 7x + 1 x2 + 7x + 13

jimthompson5910 (jim_thompson5910):

f(x) – g(x) = (2x^2 – 2x + 7) - (x^2 + 9x – 6) f(x) – g(x) = 2x^2 – 2x + 7 - x^2 - 9x + 6 f(x) – g(x) = ???

OpenStudy (anonymous):

C?

jimthompson5910 (jim_thompson5910):

2x^2 - x^2 is what

jimthompson5910 (jim_thompson5910):

this is the part where you jump in dreamer05 (don't worry if you're wrong)

OpenStudy (rulnick):

f(x)-g(x) = (2x2 -2x + 7) - (x2 + 9x - 6) = (2x2 - x2) + (-2x - 9x) + (7 + 6) = x2 - ... (do you see it?)

OpenStudy (anonymous):

Well i was thinking. 3x^2 but I know thats wrong..

jimthompson5910 (jim_thompson5910):

2x^2 -x^2 is really the same as 2x^2 - 1x^2

OpenStudy (anonymous):

So x^2 ?

jimthompson5910 (jim_thompson5910):

to combine 2x^2 - 1x^2 you subtract the coefficients so.... 2-1 = 1 which means 2x^2 - 1x^2 = 1x^2 or just x^2

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

So its not C?

jimthompson5910 (jim_thompson5910):

for the x terms -2x - 9x = -11x because -2 - 9 = -11 (all you focus on are the coefficients)

OpenStudy (anonymous):

B? Makes since to me.

jimthompson5910 (jim_thompson5910):

you have to look at the whole picture

jimthompson5910 (jim_thompson5910):

the constants are 7 and 6, they add to ???

OpenStudy (anonymous):

13

jimthompson5910 (jim_thompson5910):

so to sum things up f(x) – g(x) = (2x^2 – 2x + 7) - (x^2 + 9x – 6) f(x) – g(x) = 2x^2 – 2x + 7 - x^2 - 9x + 6 f(x) – g(x) = (2x^2 - 1x^2) + (-2x - 9x) + (7 + 6) f(x) – g(x) = (1x^2) + (-11x) + (13) f(x) – g(x) = x^2 - 11x + 13

OpenStudy (anonymous):

Okay I seee. Thank you.

jimthompson5910 (jim_thompson5910):

yw